# Qipeng Kuang Ondřej Kuželka Yuanhong Wang Yuyi Wang

### v1

@@ 3.3 Applications
In this section, we show how WCP can be used for computing WFOMC of extended $\text{C}^{2}$ sentences with several combinatorial axioms in polynomial time, i.e. we show that the fragments together with each of these axioms are domain-liftable. Some of these axioms were previously shown to be domain-liftable (e.g., connectedness axioms [MBS23], tree axiom [vBK23] and forest axiom [MBS23]), while for others we show their domain-liftability for the first time (e.g., generalized connectedness axiom and bipartite axiom). We also show that WCP allows us to efficiently compute Tutte polynomial of any graph that can be encoded by an extended $\text{C}^{2}$ sentence, e.g., block-structured graphs.
@@ 3.3.1 Axioms: Existing, Generalized, and New
$k$-connected-component axiom
As in Proposition 3.1, WCP associates the weight of each model with a term indicating the number of connected components in $G(R)$, therefore we can restrict the number of connected components of $G(R)$. The $k$-connected-component axiom to the binary predicate $R$, denoted by $connected_{k}(R)$, requires $R$ to be symmetric and $G(R)$ to consist of exactly $k$ connected components. As a special case, $connected_{1}(R)$, which represents that $G(R)$ is an undirected connected graph, i.e., has just one connected component, was already shown to be domain-liftable in [MBS23]. Here we generalize their result to $connected_{k}(R)$.
@@
The fragment of extended $\text{C}^{2}$ with a single $k$-connected-component axiom is domain-liftable.
@@
Consider an extended $\text{C}^{2}$ sentence $\Psi$, a domain of size $n$, the weighting functions $w,\overline{w}$, and a distinguished binary predicate $R$. By Proposition 3.1, we have that $\mathsf{WFOMC}(\Psi\land connected_{k}(R),n,w,\overline{w})$ is equal to $[u^{k}]f_{n}(u-1;\Psi_{R},w,\overline{w},R)$. ∎
Bipartite axiom
The bipartite axiom to the binary predicate $R$, denoted by $bipartite(R)$, requires $R$ to be symmetric and $G(R)$ to be a bipartite graph777 Note that bipartite graphs are not finitely axiomatizable in first-order logic by the compactness theorem..
@@
The fragment of extended $\text{C}^{2}$ with a single bipartite axiom is domain-liftable.
@@
Consider an extended $\text{C}^{2}$ sentence $\Psi$, a domain of size $n$ weighting functions $w,\overline{w}$, and a distinguished binary predicate $R$. We need another two predicates to represent the bipartite graph. Define
@@
$\land\forall x\forall y\ \left(((P_{1}(x)\land P_{1}(y))\lor(P_{2}(x)\land P_{2}(y)))\to\lnot R(x,y)\right),$
@@
$\mathsf{WFOMC}(\Psi\land bipartite(R),n,w,\overline{w})=\sum_{\mu\in\mathcal{M}_{\Psi_{b},n}}W(\mu,w,\overline{w})\cdot(\frac{1}{2})^{cc(\mu_{R})}=f_{n}(-\frac{1}{2};\Psi_{b},w,\overline{w},R).$
Tree axiom
The tree axiom [vBK23] to the binary predicate $R$ (denoted by $tree(R)$) requires $R$ to be symmetric and $G(R)$ to be a tree. To handle the tree axiom, we extend the ability of WCP to classify the models of a sentence by both the number of connected components in $G(R)$ and the number of edges in $G(R)$ by introducing a new variable to the weight of $R$.
@@
The fragment of extended $\text{C}^{2}$ with a single tree axiom is domain-liftable.
@@
Consider an extended $\text{C}^{2}$ sentence $\Psi$, a domain of size $n$, weighting functions $w,\overline{w}$ and a distinguished binary predicate $R$. By definition, the tree axiom requires that $G(R)$ should be connected and the number of edges in $G(R)$ should be $n-1$. By Corollary 3.1, we have
@@
Forest axiom
The forest axiom [MBS23] to the binary predicate $R$ (denoted by $forest(R)$) requires $R$ to be symmetric and $G(R)$ to be a forest. Similar to the tree axiom, we can also represent that the forest axiom is domain-liftable by Corollary 3.1.
@@
The fragment of extended $\text{C}^{2}$ with a single forest axiom is domain-liftable.
@@
Consider an extended $\text{C}^{2}$ sentence $\Psi$, a domain of size $n$, weighting functions $w,\overline{w}$ and a distinguished binary predicate $R$. By definition, the forest axiom requires the number of edges in $G(R)$ to be equal to the difference between the number of vertices in $G(R)$ and the number of connected components in $G(R)$. By Corollary 3.1, we have
@@
The tree axiom and the forest axiom have already been shown to be domain-liftable in [vBK23] and [MBS23] respectively. Nevertheless, our method provides a more flexible command of the graph structures. For example, it is easy to add further constraints on the number of trees in the forest by our method.
The strong expressive power of extended $\text{C}^{2}$ fragment and the axioms above allows us to represent many interesting combinatorial problems as well as Statistical Relational Learning (SRL) problems that can be solved in polynomial time.
@@
Consider a permutation $p_{1},\dots,p_{n}$ over $[n]$, where $p_{i}$ is the position of $i$ in the permutation. The permutation can be regarded as a directed graph with several disjoint cycles if we add an edge from $i$ to $p_{i}$ for each $i\in[n]$. Counting the number of permutations with exactly $k$ cycles is a well-known combinatorial problem whose solution is given by the unsigned Stirling numbers of the first kind. This problem can be readily solved by the $k$-connected-component axiom, where we allow the binary relation $R$ to be asymmetric. More precisely, we add the $k$-cycle-permutation axiom to the binary predicate $R$, denoted by $perm_{k}(R)$, which requires $G(R)$ to be the graph corresponding to a permutation with exactly $k$ cycles. Given an extended $\text{C}^{2}$ sentence $\Psi$, a domain of size $n$, weighting functions $w,\overline{w}$ and a distinguished binary predicate $R$, $\mathsf{WFOMC}(\Psi\land perm_{k}(R),n,w,\overline{w})$ is domain-liftable. In fact, let
@@
For SRL problems, we can express hard constraints on some relation by adding axioms to the relation, e.g., expressing a company hierarchy with the tree axiom, and player matching in a game with the bipartite axiom. In the following example, we show that more complex soft constraints can also be expressed with WCP.
@@ 3.3.2 From WCP to Tutte Polynomial
We can further use the extended WCP in Definition 3.2 to obtain the well-known Tutte polynomial of graphs that can be encoded by a fixed $\text{C}^{2}$ sentence with cardinality constraints and unary evidence. Note that the size of the $\text{C}^{2}$ sentence should be fixed to a constant independent of the graph size while the parameters for cardinality constraints and the size of unary evidence are unbounded.
The weighting functions are all set to $\mathds{1}$ (i.e., a function always returns $1$) in the rest of this section.
@@
Let $S_{1},S_{2},\cdots$ be an infinite sequence of sets of undirected graphs where each graph in $S_{n}$ has $n$ vertices. Suppose that there is a fixed $\text{C}^{2}$ sentence $\Psi$ with a distinguished symmetric binary relation $E$ such that for each $n$ and each $G$ in $S_{n}$, there is an extended $\text{C}^{2}$ sentence $\Psi_{G}$ obtained by $\Psi$ conjuncted with cardinality constraints and unary evidence, and $G$ is isomorphic to $G(\mu_{E})$ for each model $\mu\in\mathcal{M}_{\Psi_{G},n}$. [^9] For instance, [KSS22] gave the characterization of such graphs when the sentence is restricted to be in $\text{C}^{2}$. Note that in their work the counting parameters and the size of the sentence are unbounded while in our case they should be constant to apply the fast calculation of WCP. Then the Tutte polynomial of any $G$ in any $S_{n}$ can be computed as follows in time polynomial in $n$:
@@
By Proposition 3.3, computing $f_{n}\left((x-1)(y-1)-1,\sqrt{y-1};\Psi_{T}(G),\mathds{1},\mathds{1},R\right)$ is in polynomial time in $n$. $\mathsf{WFOMC}(\Psi_{G},n,\mathds{1},\mathds{1})$ is domain-liftable by the algorithm described in Section 2.4 as well. Therefore, we can obtain $T(x,y)$ in time polynomial in $n$. ∎
@@
We can get the Tutte polynomial for a simple complete graph of any size $n$. Let $S_{n}$ be the set consisting the only complete graph $G_{n}$ of $n$ vertices and define
@@
$\Psi_{G_{n}}$ $=\Psi\land|E|=n(n-1)$
for each $G_{n}$.
Then $\Psi_{G_{n}}$ consists a unique model $\mu$ over the domain of size $n$ such that $G(\mu_{E})$ is identical to $G_{n}$. Therefore,
$\frac{f_{n}\left((x-1)(y-1)-1,\sqrt{y-1};\Psi_{T}(G_{n}),\mathds{1},\mathds{1},R\right)}{(x-1)(y-1)^{n}}$
@@
Consider the Tutte polynomial of a block-structured graph defined in Example 2.4. We first fix the number $k$ of blocks and the connection among blocks. Let $S_{n}$ be the set of block-structured graphs of $k$ blocks and $n$ vertices and define similarly to Equation 2:
@@
$\Psi_{G_{n}}=$ $\Psi\land\bigwedge_{i=1}^{n}\left(Block_{v(i)}(i)\land\bigwedge_{j\neq v(i)}\lnot Block_{j}(i)\right)$
for each $G_{n}$ in each $S_{n}$. Then, the Tutte polynomial of $G_{n}$ is
$\frac{f_{n}\left((x-1)(y-1)-1,\sqrt{y-1};\Psi_{T}(G_{n}),\mathds{1},\mathds{1},R\right)}{(x-1)^{cc(G_{n})}(y-1)^{n}}.$
By Theorem 3.5, the Tutte polynomial of $G_{n}$ can be computed in polynomial time in $n$. Applying the same proof to any constant $k$ and any possible connection among blocks, we conclude that the Tutte polynomial of any block-structured graph can be computed in polynomial time in the size of the graph.
The above result answers the question in [vBK23] whether calculating the Tutte polynomial of a block-structured graph is hard. However, we remark that this result cannot be extended to computing Tutte polynomial or evaluating any point for a general graph as representing a general graph by a first-order sentence involves a list of binary evidence (i.e., a set of ground literals of arity $2$), making the WFOMC task #$\mathsf{\#P}$-hard [dBD12]. Therefore we are consistent with the hardness of Tutte polynomial.
@@ 4.1 Definition
Let $\Psi$ be a first-order logic sentence possibly with cardinality constraints and unary evidence, $w$ and $\overline{w}$ be two weighting functions and $R$ be a distinguished binary relation. The $n$-th Strict Strong Connectedness Polynomial (SSCP) $g_{n}(u,v)$ of $\Psi$ for the relation $R$ is the unique bivariate polynomial which satisfies:
@@
$\wedge$ $\forall x\forall y\left(R(x,y)\to\lnot B_{v}(x)\right)$
@@
By Corollary 2.1 and Corollary 2.2, $\chi_{G(\mu_{R})}(v)$ and $\bar{\chi}_{G(\mu_{R})}(v)$ are polynomials of $v$ which has degree at most $n$. Since $cc(\mu_{R})\leq n$, the proposition follows immediately from Proposition 4.1. ∎
@@ 4.2 Fast Calculation
Similarly to WCP, we can efficiently compute both SCPs for extended $\text{C}^{2}$ sentences. We first elaborate the algorithm for NSCP, from which the algorithm for SSCP can be easily extended.
The algorithm, also based on DP, resembles the fast calculation of WCP with the distinction that WCP involves the 1-type structure of $u+1$ layers while here we are dealing with a $(u+1)\times(v+1)$ grid. With the technique reducing the WFOMC of an extended $\text{C}^{2}$ sentence to the WFOMC of an $\text{FO}^{2}$ sentence (refer to Section 2.4), in what follows, we assume that the input sentence is an $\text{FO}^{2}$ sentence.
By Proposition 4.2, $(n+1)^{2}$ point evaluations are sufficient to interpolate the polynomial (e.g., by 2D Lagrange Interpolation). Therefore, our task is to evaluate NSCP at $u=0,1,\cdots,n$ and $v=0,1,\cdots,n$. Let $\Psi$ and $R$ be the $\text{FO}^{2}$ sentence and the binary relation involved in the NSCP. Consider the point evaluation of the NSCP at the point $(u,v)$.
@@
Observation 4.1.
Reformulate the WFOMC of $\Psi_{R,u,v}$ to the WFOMC of the sentence $\forall x\forall y\ \varphi(x,y)$, where $\varphi(x,y)$ is a quantifier-free formula. Let $\varphi_{(i,i^{\prime},i^{\prime\prime}),(j,j^{\prime},j^{\prime\prime})}(x,y)$ be the simplification of $\varphi(x,y)\land\varphi(y,x)$ when $x$ and $y$ belong to 1-type $(i,i^{\prime},i^{\prime\prime})$ and $(j,j^{\prime},j^{\prime\prime})$ respectively. Define
$r_{(i,i^{\prime},i^{\prime\prime}),(j,j^{\prime},j^{\prime\prime})}=\mathsf{WMC}(\varphi_{(i,i^{\prime},i^{\prime\prime}),(j,j^{\prime},j^{\prime\prime})}(a,b),w,\overline{w}).$
It holds that
$r_{(i,i^{\prime},i^{\prime\prime}),(j,j^{\prime},j^{\prime\prime})}=\begin{cases}r_{i,j}=\mathsf{WMC}(\psi_{i,j}(a,b),w,\overline{w}),&i^{\prime}=j^{\prime}\land i^{\prime\prime}=j^{\prime\prime},\\ r^{\to}_{i,j}=\mathsf{WMC}(\psi_{i,j}(a,b)\wedge\lnot R(b,a),w,\overline{w}),&i^{\prime}=j^{\prime}\land i^{\prime\prime}\neq j^{\prime\prime},\\ \begin{aligned} r^{\not-}_{i,j}=\mathsf{WMC}(&\psi_{i,j}(a,b)\wedge\lnot R(a,b)\wedge\lnot R(b,a),w,\overline{w}),\end{aligned}&i^{\prime}\neq j^{\prime},\\ \end{cases}$
@@
Figure 2: The 1-type structure of $\Psi_{R,u,v}$ which is a $(v+1)\times(u+1)$ grid. Each tiny gray box represents an original 1-type of $\Psi$. The bold black cells refer to the replicas of the original 1-types. The DP algorithm proceeds first along the vertical direction (from orange to blue cells) and then along the horizontal direction (from red to orange and blue cells).
@@
Input: Sentence $\Psi$, weighting functions $w,\overline{w}$, the relation $R$, domain size $n$
Output: $g_{n}(u,v;\Psi,w,\overline{w},R)$ if $w^{|}_{in}(\bm{c^{*}})$ in Algorithm 2 is given by Equation 5, or $\bar{g}_{n}(u,v;\Psi,w,\overline{w},R)$ if it is given by Equation 6
@@
Computing NSCP for a given extended $\text{C}^{2}$ sentence, a domain of size $n$, a distinguished binary relation, and weighting functions can be completed in polynomial time in $n$.
@@
Computing SSCP for a given extended $\text{C}^{2}$ sentence, a domain of size $n$, a distinguished binary relation, and weighting functions can be completed in polynomial time in $n$.
@@
There are two differences between computing SSCP and computing NSCP. The first is that the definition of SSCP implies $\forall x\lnot R(x,x)$ implicitly. The second is that $R(a,b)$ and $R(b,a)$ should be false if $a,b$ are in the 1-types in the same cell, i.e., $r_{(i,i^{\prime},i^{\prime\prime}),(j,j^{\prime},j^{\prime\prime})}=r^{\not-}_{i,j}$ when $i^{\prime}=j^{\prime}\land i^{\prime\prime}=j^{\prime\prime}$. Hence, a simple adaptation to Equation 5 that removes 1-types containing $R(x,x)$ and replaces $r_{i,j}$ with $r^{\not-}_{i,j}$:
$w^{|}_{in}(\bm{c^{*}})=\binom{|\bm{c}^{*}|}{\bm{c}^{*}}\prod_{\begin{subarray}{c}i=1,\\ \lnot R(x,x)\in C_{i}\end{subarray}}^{L}W(C_{i},w,\overline{w})^{c^{*}_{i}}\cdot s_{i}^{\binom{c^{*}_{i}}{2}}\prod_{\begin{subarray}{c}1\leq i<j\leq L,\\ \lnot R(x,x)\in C_{i},\lnot R(x,x)\in C_{j}\end{subarray}}\left(r^{\not-}_{i,j}\right)^{c^{*}_{i}\cdot c^{*}_{j}},$ (6)

### v2

@@ 3.3 Applications
In this section, we show how WCP can be used for computing WFOMC of $\text{C}^{2}$ sentences with cardinality constraints and several combinatorial axioms in polynomial time, i.e. we show that the fragments together with each of these axioms are domain-liftable. Some of these axioms were previously shown to be domain-liftable (e.g., connectedness axioms [MBS23], tree axiom [vBK23] and forest axiom [MBS23]), while for others we show their domain-liftability for the first time (e.g., generalized connectedness axiom and bipartite axiom). We also show that WCP allows us to efficiently compute the Tutte polynomial of any graph that can be encoded by a set of ground unary literals, cardinality constraints and a fixed $\text{C}^{2}$ sentence, e.g., block-structured graphs.
@@ 3.3.1 Axioms: Existing, Generalized, and New
$k$-connected-component axiom. As in Proposition 3.2, WCP associates the weight of each model with a term indicating the number of connected components in $G(R)$, therefore we can restrict the number of connected components of $G(R)$. The $k$-connected-component axiom to the binary predicate $R$, denoted by $connected_{k}(R)$, requires $R$ to be symmetric and $G(R)$ to consist of exactly $k$ connected components. As a special case, $connected_{1}(R)$, which represents that $G(R)$ is an undirected connected graph, i.e., has just one connected component, was already shown to be domain-liftable in [MBS23]. Here we generalize their result to $connected_{k}(R)$.
@@
The fragment of $\text{C}^{2}$ with cardinality constraints and a single $k$-connected-component axiom is domain-liftable.
@@
Consider a $\text{C}^{2}$ sentence $\Psi$ with cardinality constraints, a domain of size $n$, the weighting functions $w,\overline{w}$, and a distinguished binary predicate $R$. By Proposition 3.2, we have that $\mathsf{WFOMC}(\Psi\land connected_{k}(R),n,w,\overline{w})$ is equal to $[u^{k}]f_{n}(u-1;\Psi_{R},w,\overline{w},R)$. ∎
Bipartite axiom. The bipartite axiom to the binary predicate $R$, denoted by $bipartite(R)$, requires $R$ to be symmetric and $G(R)$ to be a bipartite graph888 Note that bipartite graphs are not finitely axiomatizable in first-order logic by the compactness theorem..
@@
The fragment of $\text{C}^{2}$ with cardinality constraints and a single bipartite axiom is domain-liftable.
@@
Consider a $\text{C}^{2}$ sentence $\Psi$ with cardinality constraints, a domain of size $n$ weighting functions $w,\overline{w}$, and a distinguished binary predicate $R$. We need another two predicates to represent the bipartite graph. Define
@@
$\land\forall x\forall y\ \left(P_{1}(x)\land P_{1}(y)\lor P_{2}(x)\land P_{2}(y)\to\lnot R(x,y)\right),$
@@
$\mathsf{WFOMC}(\Psi\land bipartite(R),n,w,\overline{w})$ $=\sum_{\mu\in\mathcal{M}_{\Psi_{b},n}}W(\mu,w,\overline{w})\cdot(\frac{1}{2})^{cc(\mu_{R})}$
$=f_{n}(-\frac{1}{2};\Psi_{b},w,\overline{w},R).$
Tree axiom. The tree axiom [vBK23] to the binary predicate $R$ (denoted by $tree(R)$) requires $R$ to be symmetric and $G(R)$ to be a tree. To handle the tree axiom, we extend the ability of WCP to classify the models of a sentence by both the number of connected components in $G(R)$ and the number of edges in $G(R)$ by introducing a new variable to the weight of $R$.
@@
The fragment of $\text{C}^{2}$ with cardinality constraints and a single tree axiom is domain-liftable.
@@
Consider a $\text{C}^{2}$ sentence $\Psi$ with cardinality constraints, a domain of size $n$, weighting functions $w,\overline{w}$ and a distinguished binary predicate $R$. By definition, the tree axiom requires that $G(R)$ should be connected and the number of edges in $G(R)$ should be $n-1$. By Corollary 3.9, we have
@@
Forest axiom. The forest axiom [MBS23] to the binary predicate $R$ (denoted by $forest(R)$) requires $R$ to be symmetric and $G(R)$ to be a forest. Similar to the tree axiom, we can also represent that the forest axiom is domain-liftable by Corollary 3.9.
@@
The fragment of $\text{C}^{2}$ with cardinality constraints and a single forest axiom is domain-liftable.
@@
Consider a $\text{C}^{2}$ sentence $\Psi$ with cardinality constraints, a domain of size $n$, weighting functions $w,\overline{w}$ and a distinguished binary predicate $R$. By definition, the forest axiom requires the number of edges in $G(R)$ to be equal to the difference between the number of vertices in $G(R)$ and the number of connected components in $G(R)$. By Corollary 3.9, we have
@@
The tree axiom and the forest axiom have already been shown to be domain-liftable in [vBK23] and [MBS23] respectively. Nevertheless, our method provides a more flexible command of the graph structures. For example, it is easy to add further constraints on the number of trees in the forest by our method (we will discuss this in Section 5).
The strong expressive power of $\text{C}^{2}$ fragment with cardinality constraints and the axioms above allows us to represent many interesting combinatorial problems as well as Statistical Relational Learning (SRL) problems.
@@
Consider a permutation $p_{1},\dots,p_{n}$ over $[n]$, where $p_{i}$ is the position of $i$ in the permutation. The permutation can be regarded as a directed graph with several disjoint cycles if we add an edge from $i$ to $p_{i}$ for each $i\in[n]$. Counting the number of permutations with exactly $k$ cycles is a well-known combinatorial problem whose solution is given by the unsigned Stirling numbers of the first kind. This problem can be readily solved by the $k$-connected-component axiom, where we allow the binary relation $R$ to be asymmetric. More precisely, we add the $k$-cycle-permutation axiom to the binary predicate $R$, denoted by $perm_{k}(R)$, which requires $G(R)$ to be the graph corresponding to a permutation with exactly $k$ cycles. Given a $\text{C}^{2}$ sentence $\Psi$ with cardinality constraints, a domain of size $n$, weighting functions $w,\overline{w}$ and a distinguished binary predicate $R$, $\mathsf{WFOMC}(\Psi\land perm_{k}(R),n,w,\overline{w})$ is domain-liftable. In fact, let
@@
In the field of SRL, we can express hard constraints on some relation by adding axioms to the relation, e.g., expressing a company hierarchy with the tree axiom, and player matching in a game with the bipartite axiom. In the following example, we show that more complex soft constraints can also be expressed with WCP.
@@ 3.3.2 From WCP to Tutte Polynomial
We can further use the extended WCP in Definition 3.8 to obtain the well-known Tutte polynomial of undirected graphs that can be encoded by a set of ground unary literals, cardinality constraints and a fixed $\text{C}^{2}$ sentence. Note that the size of the $\text{C}^{2}$ sentence should be fixed to a constant independent of the graph size (i.e., the number of vertices) while the parameters for cardinality constraints and the size of ground unary literals are unbounded. The weighting functions are all set to $\mathds{1}$ (i.e., a function always returns $1$) in the rest of this section.
@@
Let $S$ be an infinite set of undirected graphs. Suppose that there is a fixed $\text{C}^{2}$ sentence $\Psi$ with a distinguished symmetric binary relation $E$ such that for each graph $G\in S$ of size $n$, there is a $\text{C}^{2}$ sentence $\Psi_{G}$ obtained by conjuncting $\Psi$ with cardinality constraints and ground unary literals, and $G$ is isomorphic to $G(\mu_{E})$ for each model $\mu\in\mathcal{M}_{\Psi_{G},n}$. [^10] For instance, [KSS22] gave the characterization of such graphs when the sentence is restricted to be in $\text{C}^{2}$. Note that in their work the counting parameters and the size of the sentence are unbounded while in our case they should be constant to apply the fast calculation of WCP. Then the Tutte polynomial of any $G\in S$ of size $n$ can be computed as follows in time polynomial in $n$:
@@
By Proposition 3.5, computing $f_{n}\left((x-1)(y-1)-1,\sqrt{y-1};\Psi_{T}(G),\mathds{1},\mathds{1},R\right)$ is in polynomial time in $n$. Since the cardinality of $S$ is infinite, the size of $\Psi$ is independent of the size of $G$, thus $\mathsf{WFOMC}(\Psi_{G},n,\mathds{1},\mathds{1})$ is domain-liftable by the algorithm described in Section 2.4 as well. Therefore, we can obtain $T(x,y)$ in time polynomial in $n$. ∎
@@
We can get the Tutte polynomial for a simple undirected complete graph of any size $n$. Let $S$ be the set of all simple undirected complete graphs and define
@@
$\Psi_{G}$ $=\Psi\land(|E|=n(n-1))$
for each $G\in S$ of size $n$.
Then $\Psi_{G}$ has a unique model $\mu$ over the domain of size $n$ such that $G(\mu_{E})$ is identical to $G$. Therefore,
$\frac{f_{n}\left((x-1)(y-1)-1,\sqrt{y-1};\Psi_{T}(G),\mathds{1},\mathds{1},R\right)}{(x-1)(y-1)^{n}}$
@@
Consider the Tutte polynomial of a block-structured graph defined in Example 2.10. We first fix the number $k$ of blocks and the connection among blocks. Let $S$ be the set of block-structured graphs of $n$ vertices such that each graph has $k$ blocks and the fixed connection. Define similarly to Equation (2) the following sentences:
@@
$\Psi_{G}=$ $\Psi\land\bigwedge_{e=1}^{n}\left(Block_{v(e)}(e)\land\bigwedge_{j\neq v(e)}\lnot Block_{j}(e)\right)$
for each $G\in S$ of size $n$. Then, the Tutte polynomial of $G$ is
$\frac{f_{n}\left((x-1)(y-1)-1,\sqrt{y-1};\Psi_{T}(G),\mathds{1},\mathds{1},R\right)}{(x-1)^{cc(G)}(y-1)^{n}}.$
By Theorem 3.15, the Tutte polynomial of $G$ can be computed in polynomial time in $n$. Applying the same proof to any constant $k$ and any possible connection among blocks, we conclude that the Tutte polynomial of any block-structured graph can be computed in polynomial time in the size of the graph.
The above result answers positively the question in [vBK23] whether calculating the Tutte polynomial of a block-structured graph is hard. However, we remark that this result cannot be extended to computing the Tutte polynomial or evaluating any point for general graphs since representing a general graph by a first-order sentence involves a list of ground binary literals, making the WFOMC task #$\mathsf{\#P}$-hard [dBD12]. Therefore we are consistent with the hardness of computing the Tutte polynomial of general graphs [JVW90], while showing tractability for block-structured ones (with a constant number of blocks).
@@ 4.1 Definition
Let $\Psi$ be a first-order logic sentence possibly with cardinality constraints and ground unary literals, $w$ and $\overline{w}$ be two weighting functions and $R$ be a distinguished binary relation. The $n$-th Strict Strong Connectedness Polynomial (SSCP) $g_{n}(u,v)$ of $\Psi$ for the relation $R$ is the unique bivariate polynomial which satisfies:
@@
$\wedge$ $\forall x\forall y\left(R(x,y)\to\lnot B_{v}(x)\land B_{1}(y)\right)$
@@
By Corollary 2.14 and 2.15, $\chi_{G(\mu_{R})}(v)$ and $\bar{\chi}_{G(\mu_{R})}(v)$ are polynomials of $v$ which has degree at most $n$. Since $cc(\mu_{R})\leq n$, the proposition follows immediately from Proposition 4.2. ∎
@@ 4.2 Fast Calculation
Similarly to WCP, we can efficiently compute both SCPs for $\text{C}^{2}$ sentences with cardinality constraints. We first elaborate the algorithm for NSCP, from which the algorithm for SSCP can be easily extended.
The algorithm, also based on DP, resembles the fast calculation of WCP with the distinction that WCP involves the 1-type structure of $u+1$ layers while here we are dealing with a $(u+1)\times(v+1)$ grid. With the technique reducing WFOMC of $\text{C}^{2}$ with cardinality constriants to WFOMC of $\text{FO}^{2}$ (refer to Section 2.4), in what follows, we assume that the input sentence is an $\text{FO}^{2}$ sentence.
By Proposition 4.3, $(n+1)^{2}$ point evaluations at $u=0,1,\cdots,n$ and $v=0,1,\cdots,n$ are sufficient to interpolate the polynomial (e.g., by 2D Lagrange Interpolation). Therefore, our task is to evaluate NSCP at these points. Let $\Psi$ and $R$ be the $\text{FO}^{2}$ sentence and the binary relation involved in the NSCP. Consider the point evaluation of the NSCP at the point $(u,v)$.
@@
Observation 4.4.
Let $r_{(i,i^{\prime},i^{\prime\prime}),(j,j^{\prime},j^{\prime\prime})}$ be the mutual 1-type coefficient (defined in Section 2.4) of $\Psi_{R,u,v}$. It holds that
$r_{(i,i^{\prime},i^{\prime\prime}),(j,j^{\prime},j^{\prime\prime})}=\begin{cases}r_{i,j}=\mathsf{WMC}(\psi_{i,j}(a,b),w,\overline{w}),&i^{\prime}=j^{\prime}\land i^{\prime\prime}=j^{\prime\prime},\\ r^{\to}_{i,j}=\mathsf{WMC}(\psi_{i,j}(a,b)\wedge\lnot R(b,a),w,\overline{w}),&i^{\prime}=j^{\prime}\land i^{\prime\prime}\neq j^{\prime\prime},\\ \begin{aligned} r^{\not-}_{i,j}=\mathsf{WMC}(&\psi_{i,j}(a,b)\wedge\lnot R(a,b)\\ &\wedge\lnot R(b,a),w,\overline{w}),\end{aligned}&i^{\prime}\neq j^{\prime},\\ \end{cases}$
@@
Figure 2: The 1-type structure of $\Psi_{R,u,v}$ which is a $(v+1)\times(u+1)$ grid. Each tiny gray box represents an original 1-type of $\Psi$. The bold gray cells (each containing $4\times 4$ boxes) refer to the replicas of the original 1-types. The DP algorithm proceeds first along the vertical direction (from cell $(4,0)$ to cell $(4,4)$) and then along the horizontal direction (from columns $0,\cdots,3$ to column $4$).
@@
Input: A sentence $\Psi$, weighting functions $w,\overline{w}$, the relation $R$, and an integer $n$
Output: $g_{n}(u,v;\Psi,w,\overline{w},R)$ if $w^{|}_{in}(\bm{c^{*}})$ in Algorithm 2 is given by Equation (5), or $\bar{g}_{n}(u,v;\Psi,w,\overline{w},R)$ if it is given by Equation (6)
@@
Computing NSCP for a given $\text{C}^{2}$ sentence with cardinality constraints, a domain of size $n$, a distinguished binary relation, and weighting functions can be completed in polynomial time in $n$.
@@
Computing SSCP for a given $\text{C}^{2}$ sentence with cardinality constraints, a domain of size $n$, a distinguished binary relation, and weighting functions can be completed in polynomial time in $n$.
@@
There are two differences between computing SSCP and computing NSCP. The first is that the definition of SSCP implies $\forall x\lnot R(x,x)$ implicitly. The second is that $R(a,b)$ and $R(b,a)$ should be false if $a,b$ are in the 1-types in the same cell, i.e., $r_{(i,i^{\prime},i^{\prime\prime}),(j,j^{\prime},j^{\prime\prime})}=r^{\not-}_{i,j}$ when $i^{\prime}=j^{\prime}\land i^{\prime\prime}=j^{\prime\prime}$ and $s_{i}$ should be replaced with
$s^{\not-}_{i}=\mathsf{WMC}(\psi_{i}(a,b)\land\psi_{i}(b,a)\land\lnot R(a,b)\land\lnot R(b,a),w,\overline{w}).$
Hence, a simple adaptation to Equation (5) that removes 1-types containing $R(x,x)$ and replaces $r_{i,j}$ and $s_{i}$ with $r^{\not-}_{i,j}$ and $s^{\not-}_{i}$:
$w^{|}_{in}(\bm{c^{*}})=\binom{|\bm{c}^{*}|}{\bm{c}^{*}}\prod_{\begin{subarray}{c}i=1,\\ \lnot R(x,x)\in C_{i}\end{subarray}}^{L}W(C_{i},w,\overline{w})^{c^{*}_{i}}\cdot\left(s^{\not-}_{i}\right)^{\binom{c^{*}_{i}}{2}}\prod_{\begin{subarray}{c}1\leq i<j\leq L,\\ \lnot R(x,x)\in C_{i},\lnot R(x,x)\in C_{j}\end{subarray}}\left(r^{\not-}_{i,j}\right)^{c^{*}_{i}\cdot c^{*}_{j}}$ (6)